# On Meta-Prompting

### v2

@@ 4.2.1. Agnosticity
Mathematically, to prove this we require showing the converse, in a sense, to Lemma 1: there will always exist a set of meta-prompt morphisms that can take an arbitrary task description as an input, and return relevant outputs, even if there does not exist a functor between these tasks. We prove this in Theorem 1 below. Intuitively, it is a consequence from the fact that Prompt is right-closed, and by definition it has an internal hom and an isomorphism $\text{Hom}_{\text{{Prompt}}}(X\otimes Y,Z)\simeq\text{Hom}_{\text{{Prompt}}}(Y,Z^{X})$.
Theorem 1 (Task-Agnosticity of Meta-Prompt Morphisms).
@@
Proof.
(sketch; full proof in Appendix B) Recall the definition of a task category: $T_{1}\colon\text{{Task}}_{1}\xhookrightarrow{}\text{{Prompt}}{}$ (r. $T_{2},\text{{Task}}_{2}$). Hence there exists a map from the task-category’s meta-prompt morphisms and objects into Prompt. From Lemma 1 we know that the internal hom $\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}\rightarrow\text{{Prompt}}{}$ and its isomorphisms capture the system prompts (i.e., $\tilde{T_{1}},\tilde{T_{2}}$) plus user contexts as part of its construction. In other words, the internal hom encodes precisely (the hom set of) the meta-prompt morphisms for both tasks. ∎
Theorem 1 shows that one can build a general-purpose meta-prompt morphism that generates prompts by simply encoding the task description as part of the prompt, even when the tasks are not related.
4.2.2. Equivalence
A natural consequence of Theorem 1 is that all meta-prompt morphisms are equivalent in the sense that they can all be transformed between one another. We show this in Corollary 2.
Corollary 0 (Equivalence of Meta-Prompt Morphisms).
@@ 4.2.2. Equivalence
4.2.3. Implications: Comparison with Predefined System Prompts
@@
5. Experiments
We empirically evaluate our theoretical framework’s predictions in two tasks: how to improve a text, which we dub Ideation; and how to continue writing said text, or Creativity. We selected these areas based on common applications and active research areas around LLMs, and because they are not functorially related. For all tasks and data-generation aspects we utilized GPT-4 (Open AI, 2023) (version: gpt-4-0613), an instruction-pretrained model, through the Azure Open AI API. We used the default call parameters for tokens and temperature. All data analysis was done in a consumer-grade laptop. [^5] Full prompts and code to reproduce this work is in https://github.com/adewynter/metaprompting In this section, for short, we refer to the prompts (r. outputs) generated by a meta-prompt as meta-generated prompts (r. outputs).
5.1. Ideation
Ideation takes in a text and suggest ideas on how to improve it. The underlying task-category Idea has as objects passages, and as arrows ways to improve the (same) passage. Each object $X\in\text{{Idea}}$ has general-purpose arrows such as ‘‘Make it more concise’’, ‘‘Make it longer’’, and ‘‘Explain this to a 5 year old’’. These three are our baseline prompts for this task. We assume that our baseline arrows belong to every exponential object $Z^{X}\in\text{{Idea}}$, but that they also have some disjoint, context-specific arrows: for example, if the context is ‘‘doing something for a long time can be boring, having hobbies and doing sports can refresh our mind’’, an appropriate morphism would be ‘‘Add some details to support your claim that hobbies and sports can refresh our mind.’’ For our dataset we sampled 300 points from WikiText-103 (Merity et al., 2017) and the TOEFL11 corpus (Blanchard et al., 2013). The full prompt is in Appendix D.
5.2. Creativity
The underlying task-category Creat has as objects passages, and as arrows prompts that generate said linguistic-and-semantically natural transitions between them. Same as before, we assume that the exponential objects of Creat have disjoint elements. Our baseline prompts are ‘‘Write a paragraph to connect the left text and right texts,’’ ‘‘Insert a passage connecting the two passages’’, and the empty string (implicitly passage completion). For this task we sampled 300 points of TOEFL11, and version 3.0.0 of the DailyMail/CNN corpus (Hermann et al., 2015). The full prompt is in Appendix D.
5.3. Meta-Prompt
Our meta-prompt takes in a replaced string with the task description (the original system prompt) and the context. For example, if the system prompt for Creativity is ‘‘return a linguistically natural transition between {LEFT} and {RIGHT}’’, the replacement string {TASK DESCRIPTION GOES HERE} for the meta-prompt will include this description, and the context will include the passages. For illustration purposes, we show a sample meta-prompt in Prompt 1; and the full prompt in Appendix D.
Table 1. Sample meta-prompt. We replace the placeholders {IN BRACKETS} with the user content, system prompt (task description), and relevant exemplars. The anthropomorphization of the meta-prompt (‘‘you are a teacher’’) is a common technique in prompting. The rest of the instructions, along with the task-specific exemplars, are to ensure a parseable, appealing response.
# Input:
{USER CONTENT GOES HERE}
# System
You are teacher talking to a student. Propose five user commands such that {TASK DESCRIPTION GOES HERE}.
# Instructions:
- Write the commands in one sentence
- The commands should be short and concise
- Write five independent commands
# Exemplars:
1. {EXEMPLARS GO HERE}
# Commands:
1.
5.4. Data Annotation
5.5. Results
In Creativity we observed a clear separation of preferences, with both meta-generated prompts and outcomes being marked as more desirable. Meta-generated prompts were in the top 3 selections 70% of the time, and the meta-generated outputs 61% of the time. Notably, the explicit task definitions as baselines were often ranked as the least suitable suggestions (Figures 1(a) and 2(a)).
In Ideation we observed an overwhelmingly large amount of users preferring the first meta-generated prompt as opposed to the other two, as well as its relevant output (Figures 1(b) and 2(b)). Meta-generated prompts were in the top 3 selections 70% of the time, and the meta-generated outputs 61% of the time. Nonetheless, both the meta-generated prompt and its corresponding output were ranked, on average, as more desirable than the baseline prompts. The rankings are statistically significant for both tasks ($p<0.01$) under a Wilcoxon signed-rank test.
(a) Meta-generated prompts were in the top 3 selections 71% of the time. We observed a slight preference towards prompt 2.
(b) Meta-generated prompts were in the top 3 selections 70% of the time. We observed a marked skewness towards the first.
Figure 1. Rankings for prompt suitability given a context for Creativity (left) and Ideation (right). Prompts 1-3 are meta-generated prompts. Baselines are marked with an asterisk (*). Prompts 5 and 6 are the task description. Darker areas have higher frequencies: baseline prompts are ranked as least suitable.
(a) Meta-generated outputs were in the top 3 selections 59% of the time. We did not observe strong preference towards output 2, suggesting that the meta-generated outputs were evenly regarded as suitable.
(b) Meta-generated outputs were in the top 3 selections 61% of the time. For this task, we noticed a noticeable skewness in preferences towards the first prompt, which is in line with the results for prompt suitability.
Figure 2. Rankings for output suitability given a context for Creativity (left) and Ideation (right). Outputs 1-3 are generated from our meta-generated prompts. Baselines are marked with an asterisk (*). Outputs 5 and 6 are generated by calling the model with task description alone. Darker areas have higher frequencies, with baseline outputs are ranked as least suitable.
6. Discussion
@@
In our tests, the meta-generated prompts were ranked by the users as more useful than the baselines. Likewise, they were more likely to generate content regarded as more suitable. While it could be argued that, due to the model’s prompt sensitivity, a good prompt would lead to a good outcome, in Creativity we included the system prompt in our baseline prompts, and hence we had the task’s desired output as part of the outputs to be evaluated. This desired output was ranked below the meta-generated outputs.
@@
From a mathematical standpoint, Creat is a collection of morphisms that describe the task via an inclusion functor. The task description is itself a morphism in $\text{Hom}_{\text{{Creat}}}(-,-)$, and hence not sufficient to encompass the potential nuances associated with its domain (the context). That is, a (system) prompt stating ‘‘rewrite {X}’’ may not be as effective as the collection of meta-generated prompts $Z^{X}$ that, based on $X$, are able to return a contextualized phrasing, and hence a better output (e.g. ‘‘rewrite {X} in a way that the first sentence stands out more’’).
@@
7. Related Work
7.1. Prompting to Prompt
@@
Meta-prompting approaches have also found success at lowering biases (Guo et al., 2022), event argument extraction (Dong et al., 2022), and jailbreak detection (Pryzant et al., 2023). Liu et al. (2022) showed that generating knowledge automatically for question answering (QA) improves downstream performance. For reasoning, Auto-CoT (Zhang et al., 2023) propose building the pathways for reasoning in chain-of-thought prompting (Wei et al., 2022b; Kojima et al., 2022) automatically. Closer to pure meta-prompting automation, Yao et al. (2023) introduced a prompting paradigm designed for reasoning, which uses a feedback loop to generate reasoning steps. This is effective when integrating external-world grounding. LAMBADA (Kazemi et al., 2023) is a backward-chaining algorithm that focuses on reasoning by extracting propositions and performing proving via chaining. Auto-GPT (Richards, 2023) is a tool designed to iteratively solve problems by generating its own prompts and including a database for memory.
7.2. Category Theory
@@ 8.0.1. Category Theory
Likewise, our results from Corollary 2 can be strengthened. While our framework is powerful enough to make arguments and prove results such as the ones in Lemma 1; Theorem 1; and Section 4.2.3, it is quite coarse-grained and we are unable to show (for example) that one meta-prompting approach is better than another. This, however, is only a matter of equipping Prompt with extra structure and can be left for future work.
@@
8.0.2. Evaluation
We utilized only one model (GPT-4) and two tasks, because it has been claimed by the authors that this model is able to follow complex instructions (Open AI, 2023). Other LLMs may behave differently under meta-prompting, and hence our results--along with the small number of tasks--could sway the results. However, our $p$ values, and the fact that our framework does not depend on a specific LLM, suggests that our findings are extensible.
We did not evaluate reasoning scenarios, but we believe that it could be an important testing ground for meta-prompting. Indeed, it is worth noting that the recently-released Marco-o1 model (Zhao et al., 2024) is reasoning-focused and utilizes meta-prompting as a way to improve its own reasoning capabilities. That said, it is worth emphasizing that the contributions of our paper are on a theoretical framework for prompting and interaction with LLMs, since there exists sufficient empirical evidence that meta-prompting is effective in applied settings (Section 7).
There exist other approaches to prompt generation, such as information-theoretic methods (Sorensen et al., 2022), soft-prompting (Li and Liang, 2021; Lester et al., 2021), or fine-tuning approaches like LM-BFF (Gao et al., 2021). From our framework’s perspective, they are out of scope since they rely on altering the model, and one of our core assumptions is that the LLM remains fixed. That said, user experience could still be impacted by the choice of approach.
9. Conclusion
@@
We also experimentally showed meta-prompting’s better performance when compared to traditional prompting. Our results showed that user preference is strongly in favor of meta-generated prompts and outputs. We then used our framework to argue that the effectiveness of meta-prompting stems from not limiting the model’s instructions to the verbatim description of the task, but instead generating suitable, contextualized, task-related instructions.
Our work with Prompt showed that it is possible to abstract out common LLM issues such as prompt sensitivity, generalizability, and user interaction; and still be able to model LLM interaction. We plan to extend our framework into a more detailed treatment of Prompt to account for stochasticity, enrichment, and the FinVect category, and link that to user preferences. Some prompting techniques, such as chain-of-thought, can also be modeled with our framework. Hence, further work can also explore whether LLM reasoning capabilities improve under generalized meta-prompting approaches when compared to other techniques.
@@ Appendix A Proof of Lemma 1
For convenience, we restate Lemma 1 it below:
@@
Lemma 1 (Equivalence of Task-Categories).
@@
Therefore $g$ is sufficient to describe a functor $F\colon\text{{Task}}_{1}\rightarrow\text{{Task}}_{2}$. This concludes the proof. ∎
@@
Appendix B Proof of Theorem 1
We present a full proof of Theorem 1 below. A careful proof must pay special attention to the underlying mathematical abstraction (the internal hom) and its relation to the meta-prompt morphism. This is necessary because, although the inclusion into Prompt allows for a ‘‘catch all’’ argument on task-category equivalence [^7] If we say $\text{{Task}}_{1}$ relates to Prompt, and $\text{{Task}}_{2}$ relates to Prompt, intuitively, by the transitive property they both must be related., the argument does not necessary imply task agnosticity for the meta-prompt morphism, or even relatedness to such a granular level between the tasks.
@@ Appendix B Proof of Theorem 1
Theorem 1 (Task-Agnosticity of Meta-Prompt Morphisms).
@@
Hence the existence of the relevant meta-prompt morphisms is independent on the functor, and solely dependent on Prompt and $\tilde{T}_{1},\tilde{T}_{2}$. Their construction is done via the task description and inclusion functor as in Lemma 1, and by taking in the relevant context. ∎
Appendix C Proof of Corollary 2
Although in category theory it is often sufficient to prove statements with diagrams and arguing that they commute (‘‘proof by diagram’’), in this section we provide a clearer proof of Corollary 2. For convenience we restate it below, and remind the reader of the properties of the hom functor. Since Prompt is right-closed, the internal hom $\hom(-,-)$, a bifunctor, mapping objects and arrows from ${}^{\text{op}}\times\text{{Prompt}}{}$ Prompt to Prompt. In other words, for objects $(Y,A)\in\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}$ and arrows $f\colon X\rightarrow Y$, $g\colon B\rightarrow A$, for the internal hom $\text{hom}(f,g)$ the diagram below commutes
@@ Appendix C Proof of Corollary 2
Corollary 1 (Equivalence of Meta-Prompt Morphisms).
@@
Proof.
From Theorem 1 it is possible to map the categories and corresponding meta-prompt morphisms into Prompt, so we can work directly with this category instead of $\text{{Task}}_{1}$ and $\text{{Task}}_{2}$. We note that $X\otimes Y$ is an object in the category, and hence the internal hom will map:
@@
(1) $c_{x,y,z}\colon\text{Hom}_{\text{{Prompt}}}(X\otimes Y,Z)\overset{\sim}{\rightarrow}\text{Hom}_{\text{{Prompt}}{}}(Y,Z^{X}),$
@@
Appendix D Prompts Used
The prompt we used for Creativity is in Prompt 2. This task takes in two passages: left and right. One of them may be empty. The task requires the model to output a passage connecting left and right in a way that is linguistically and semantically related.
The prompt we used for Ideation is in Prompt 3. This task takes in three passages: left, content, and right. Left and right are used for grounding. The goal of Ideation is to return rewrites of the content. From a category-theoretical perspective, Idea is a restricted version of Rewrite (only creative rewrites are allowed).
The full meta-prompt used in our experiments is in Prompt 1. We enforced task agnosticity by replacing the content and its task description. This allowed us to retain our specific experimentation parameters across tasks.
Table 1. Sample meta-prompt. We replace {TASK} and {CONTEXT} with the system prompt (task description) and relevant input. The directive around no grammar or punctuation is to ensure an interesting output to test.

### v3

@@ 4.2.1 Agnosticity
Mathematically, to prove this we require showing the converse, in a sense, to Lemma 3.1: there will always exist a set of meta-prompt morphisms that can take an arbitrary task description as an input, and return relevant outputs, even if there does not exist a functor between these tasks. We prove this in Theorem 4.1 below. Intuitively, it is a consequence from the fact that Prompt is right-closed, and by definition it has an internal hom and an isomorphism $\text{Hom}_{\text{{Prompt}}}(X\otimes Y,Z)\simeq\text{Hom}_{\text{{Prompt}}}(Y,Z^{X})$.
Theorem 4.1 (Task-Agnosticity of Meta-Prompt Morphisms).
@@
Proof 4.2.
(sketch; full proof in Appendix B) Recall the definition of a task category: $T_{1}\colon\text{{Task}}_{1}\xhookrightarrow{}\text{{Prompt}}{}$ (r. $T_{2},\text{{Task}}_{2}$). Hence there exists a map from the task-category’s meta-prompt morphisms and objects into Prompt. From Lemma 3.1 we know that the internal hom $\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}\rightarrow\text{{Prompt}}{}$ and its isomorphisms capture the system prompts (i.e., $\tilde{T_{1}},\tilde{T_{2}}$) plus user contexts as part of its construction. In other words, the internal hom encodes precisely (the hom set of) the meta-prompt morphisms for both tasks.
Theorem 4.1 shows that one can build a general-purpose meta-prompt morphism that generates prompts by simply encoding the task description as part of the prompt, even when the tasks are not related.
4.2.2 Equivalence
A natural consequence of Theorem 4.1 is that all meta-prompt morphisms are equivalent in the sense that they can all be transformed between one another. We show this in .
Corollary 4.3 (Equivalence of Meta-Prompt Morphisms).
@@ 4.2.2 Equivalence
4.2.3 Implications: Comparison with Predefined System Prompts
@@
5 Discussion
@@
In our tests (Appendix D), the meta-generated prompts were ranked by the users as more useful than the baselines. Likewise, they were more likely to generate content regarded as more suitable. While it could be argued that, due to the model’s prompt sensitivity, a good prompt would lead to a good outcome, in Creativity we included the system prompt in our baseline prompts, and hence we had the task’s desired output as part of the outputs to be evaluated. This desired output was ranked below the meta-generated outputs.
@@
From a mathematical standpoint, Creat is a collection of morphisms that describe the task via an inclusion functor. The task description is itself a morphism in $\text{Hom}_{\text{{Creat}}}(-,-)$, and hence not sufficient to encompass the potential nuances associated with its domain (the context). That is, a (system) prompt stating “rewrite {X}” may not be as effective as the collection of meta-generated prompts $Z^{X}$ that, based on $X$, are able to return a contextualized phrasing, and hence a better output (e.g. “rewrite {X} in a way that the first sentence stands out more”).
@@
6 Related Work
6.1 Prompting to Prompt
@@
Meta-prompting approaches have also found success at lowering biases (Guo et al., 2022), event argument extraction (Dong et al., 2022), and jailbreak detection (Pryzant et al., 2023). Liu et al. (2022) showed that generating knowledge automatically for question answering (QA) improves downstream performance. For reasoning, Auto-CoT (Zhang et al., 2023) propose building the pathways for reasoning in chain-of-thought prompting (Wei et al., 2022c; Kojima et al., 2022) automatically. Closer to pure meta-prompting automation, Yao et al. (2023) introduced a prompting paradigm designed for reasoning, which uses a feedback loop to generate reasoning steps. This is effective when integrating external-world grounding. LAMBADA (Kazemi et al., 2023) is a backward-chaining algorithm that focuses on reasoning by extracting propositions and performing proving via chaining. Auto-GPT (Richards, 2023) is a tool designed to iteratively solve problems by generating its own prompts and including a database for memory.
6.2 Category Theory
@@ 7 Limitations
Likewise, our results from can be strengthened. While our framework is powerful enough to make arguments and prove results such as the ones in Lemma 3.1; Theorem 4.1; and Section 4.2.3, it is quite coarse-grained and we are unable to show (for example) that one meta-prompting approach is better than another. This, however, is only a matter of equipping Prompt with extra structure and can be left for future work.
@@
8 Conclusion
@@
We also experimentally showed meta-prompting’s better performance when compared to traditional prompting. We then used our framework to argue that the effectiveness of meta-prompting stems from not limiting the LLM’s instructions to the verbatim description of the task, but instead generating suitable, contextualized, task-related instructions.
Our work with Prompt showed that it is possible to abstract out common LLM ICL issues such as prompt sensitivity, generalizability, and user interaction; and still be able to model LLM interaction. We plan to extend our framework into a more detailed treatment of Prompt to account for stochasticity, enrichment, and the FinVect category, and link that to user preferences. Some prompting techniques, such as chain-of-thought, and its hybridization with Markov chains, such as Deepseek R1’s DeepSeek-AI (2025) can also be modeled with our framework, although we leave it for future work.
@@ Appendix A Proof of Lemma 1
For convenience, we restate Lemma 3.1 it below:
@@
Lemma 1 (Equivalence of Task-Categories)
@@
Therefore $g$ is sufficient to describe a functor $F\colon\text{{Task}}_{1}\rightarrow\text{{Task}}_{2}$. This concludes the proof.
@@
Appendix B Proof of Theorem 4.1
We present a full proof of Theorem 4.1 below. A careful proof must pay special attention to the underlying mathematical abstraction (the internal hom) and its relation to the meta-prompt morphism. This is necessary because, although the inclusion into Prompt allows for a “catch all” argument on task-category equivalence [^7] If we say $\text{{Task}}_{1}$ relates to Prompt, and $\text{{Task}}_{2}$ relates to Prompt, intuitively, by the transitive property they both must be related., the argument does not necessary imply task agnosticity for the meta-prompt morphism, or even relatedness to such a granular level between the tasks.
@@
Theorem 1 (Task-Agnosticity of Meta-Prompt Morphisms)
@@
Hence the existence of the relevant meta-prompt morphisms is independent on the functor, and solely dependent on Prompt and $\tilde{T}_{1},\tilde{T}_{2}$. Their construction is done via the task description and inclusion functor as in Lemma 1, and by taking in the relevant context.
Appendix C Proof of
Although in category theory it is often sufficient to prove statements with diagrams and arguing that they commute (“proof by diagram”), in this section we provide a clearer proof of . For convenience we restate it below, and remind the reader of the properties of the hom functor. Since Prompt is right-closed, the internal hom $\hom(-,-)$, a bifunctor, mapping objects and arrows from ${}^{\text{op}}\times\text{{Prompt}}{}$ Prompt to Prompt. In other words, for objects $(Y,A)\in\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}$ and arrows $f\colon X\rightarrow Y$, $g\colon B\rightarrow A$, for the internal hom $\text{hom}(f,g)$ the diagram below commutes
@@ Appendix C Proof of
Corollary 1 (Equivalence of Meta-Prompt Morphisms)
@@
Proof C.1.
From Theorem 4.1 it is possible to map the categories and corresponding meta-prompt morphisms into Prompt, so we can work directly with this category instead of $\text{{Task}}_{1}$ and $\text{{Task}}_{2}$. We note that $X\otimes Y$ is an object in the category, and hence the internal hom will map:
@@
$c_{x,y,z}\colon\text{Hom}_{\text{{Prompt}}}(X\otimes Y,Z)\overset{\sim}{\rightarrow}\text{Hom}_{\text{{Prompt}}{}}(Y,Z^{X}),$ (2)
@@
Appendix D Experiments
[…]
D.1 Meta-Prompts
Our meta-prompt takes in a replaced string with the task description (the original system prompt) and the context. For example, if the system prompt for Creativity is “return a linguistically natural transition between {LEFT} and {RIGHT}”, the replacement string {TASK DESCRIPTION GOES HERE} for the meta-prompt will include this description, and the context will include the passages. This replacement serves to illustrate task agnosticity, and allowed us to retain our specific experimentation parameters across tasks. The prompt is in Prompt 1.
D.2 Ideation
Ideation takes in a text and suggest ideas on how to improve it. The underlying task-category Idea has as objects passages, and as arrows ways to improve the (same) passage. Idea is a restricted version of Rewrite (only creative rewrites are allowed). Each object $X\in\text{{Idea}}$ has general-purpose arrows such as “Make it more concise”, “Make it longer”, and “Explain this to a 5 year old”. These three are our baseline prompts for this task. We assume that our baseline arrows belong to every exponential object $Z^{X}\in\text{{Idea}}$, but that they also have some disjoint, context-specific arrows: for example, if the context is “doing something for a long time can be boring, having hobbies and doing sports can refresh our mind”, an appropriate morphism would be “Add some details to support your claim that hobbies and sports can refresh our mind.” For our dataset we sampled 300 points from WikiText-103 (Merity et al., 2017) and the TOEFL11 corpus (Blanchard et al., 2013). The full prompt is in Prompt 2.
D.3 Creativity
The underlying task-category Creat has as objects passages, and as arrows prompts that generate said linguistic-and-semantically natural transitions between them. Same as before, we assume that the exponential objects of Creat have disjoint elements. Our baseline prompts are “Write a paragraph to connect the left text and right texts,” “Insert a passage connecting the two passages”, and the empty string (implicitly passage completion). For this task we sampled 300 points of TOEFL11, and version 3.0.0 of the DailyMail/CNN corpus (Hermann et al., 2015). The full prompt is in Prompt 3.
Table 1: Sample meta-prompt. We replace {TASK} and {CONTEXT} with the system prompt (task description) and relevant input, as described in Prompt 2 and Prompt 3. The directive around no grammar or punctuation is to ensure an interesting output to test.